5. You have collected data on the processing speed for 20 super computers, both before and after a newly designed chip was imData:

"","x"
"1",0.0458835170914729
"2",0.186357728205621
"3",0.261930461196287
"4",0.121678145570879
"5",0.190702709369361
"6",0.244160178413892
"7",1.12384053562444
"8",1.76960612885786
"9",0.360237084315125
"10",0.164134511258453
"11",0.209619600170602
"12",0.195104011489699
"13",0.163665047740577
"14",0.826605167818777
"15",0.523335928069361
"16",0.971550465234638
"17",0.664677629840183
"18",0.264826960157033
"19",0.383002542661351
"20",0.139428810965134
Answer #1

page-) Given that, collected data on the processing speed for 20 super compatess. so let us - nis sample size. - ad is mean opage 2 from the above calcalations we conclude that, confidane onrual i for the before and after Omplementation for the a offone 3 The histagram of the differance value is Positively skew, so the assumption of the normal distribution on the diffesanpage-4 - t= qt (P-o.10, df=19, low. de 19, lower fail of) [1] 1.327728 =) tq (p=orlo, df=19 power fail=f) [1] 1.327728 - 80%